# Artificial Intelligence in Creative Industries: Advances Prior to 2025

### v4

@@ 2.3 Diffusion Models
A generative model, in the context of AI, exploits machine learning to learn a probability distribution of the training data to generate new data samples. The very first models were based on Autoencoders (AEs) that learn to encode input data into a lower-dimensional representation (latent space) and then decode it back to its original form. A specific type of AE, a variational autoencoders (VAE) [31], learns the latent space as statistical parameters of probabilistic distributions, leading to significant improvement of the generated results. Concurrently, Goodfellow et al. [32] introduced an alternative architecture known as a Generative Adversarial Network (GAN). GANs comprise two competing AI modules: a generator, which creates a sample, and a discriminator, which determines whether the received sample is real or generated. When comparing VAEs to GANs, VAEs exhibit greater stability during training, whereas GANs excel at producing realistic images. More details about AEs and GANs for creative technologies can be found in our previous review [8].
@@
An important factor in driving the rapid growth of generative AI has been the development of diffusion probabilistic models (referred to as diffusion models (DMs)). The first DM was introduced in 2015 by Sohl-Dickstein et al. [33], using Nonequilibrium Thermodynamics. However, it took a further 5 years for DMs to generate desirable results: the era of DMs began with Denoising Diffusion Probabilistic Models (DDPMs) proposed by Ho et al. [34] in 2020 and Score-based diffusion models proposed by Song et al. [35] in 2021. These involve a simplified process using a denoising autoencoder to approximate Bayesian inference. In brief, the models leverage a diffusion process to learn a probability distribution of the input data. As the name suggests, the data is diffused by gradually adding noise at each iteration step as shown in Fig. 1 (b). A deep neural network (DNN) is then trained to remove this noise, called the denoising process or reverse process. Consequently, the trained model uses random noise to generate data with characteristics similar to those of the training samples. Comparing to GANs, DMs provide higher diversity samples [36] and a training process that is much more stable and does not suffer from mode collapse. DMs are however computationally intensive and require longer training times compared to GANs. The complexity can be significantly reduced by training the DMs in latent space. Latent Diffusion Models (LDM) [24] use pretrained networks to convert images to feature maps, and perform training on a low-dimensional space. The diagram of LDM is shown in Fig. 1 (c).
@@
Generating a synthesized sample at random might not be particularly useful, especially for creative industry applications. Therefore, conditional diffusion models have been proposed, supporting a wide range of applications such as text-to-sound, text-to-images, and image-to-videos. For DMs, the conditional distributions are modeled using a conditional denoising autoencoder. Classifier guidance was introduced in [36] to improve the generation of images of a desired class. For example, when we provide the model with information, such as ‘a flower’, the DM will synthesize a variety of flower images, as the word ‘flower’ guides the model toward the latent distribution that is formed by various images of flowers. The work in [37] simply refines the latent space of well-trained unconditional DDPM so that the higher-level semantics of the synthetic samples are similar to the reference (conditioning). The LDM [24] offers more flexible conditional image generators by adding cross-attention layers (referred to Transformers in Section 2.1) to the denoising autoencoder. A survey on the methods and applications of DMs prior to 2024 can be found in [38].
@@ 2.4 Implicit Neural Representations
Implicit Neural Representations (INR), also called neural fields, neural implicits or coordinate-based neural networks, represent input content implicitly through learned functions $F$, as shown in Eq. 3. They can be considered as fields $x$ (represented by a scalar, vector, or a tensor with a value, such as a magnetic field in physics) that are fully or partially parameterized by a neural network $\Phi$, typically an MLP [39].
@@
Although this concept appears complex, the process is actually very straightforward. For example, in the case of an image, the coordinates of each pixel $(x,y)$ contain color information $(r,g,b)$. The INR inputs $(x,y)$ to the MLP and learns to provide the output $(r,g,b)$. The weights and biases of the MLP now represent such an image. Usually, the number of parameters of the MLP is smaller than the total number of pixels multiplied by 3, accounting for the 3 color channels. Hence, one of its emerging applications is in data compression [40]. Moreover, the INR can handle complex and high-dimensional data efficiently, attracting attention for visual computing applications such as 3D scene reconstruction.
Traditional MLPs employ ReLU (rectified linear unit) for non-linear activation due to its simplicity. However, Sitzmann et al. [41] demonstrated that periodic functions, such as sinusoids, are more suitable for representing complex natural signals, offering a better fit to the first- and second-order derivatives of the signals. However, this activation can cause ringing artifacts. Saragadam et al. instead proposed using complex Gabor wavelets [42], which learn to represent high frequencies better and simultaneously are robust to noise.
One of the fastest-growing areas that exploits INRs is Neural Radiance Fields (NeRF), evidenced by 57 papers presented at CVPR, the largest annual conference in computer vision, in 2022 growing to 175 papers in 2023101010https://markboss.me/post/nerf_at_cvpr23/, before dropping to 71 in 2024, largely due to competition from 3D Gaussian Splatting [^11] https://github.com/Yubel426/NeRF-3DGS-at-CVPR-2024. First introduced in 2020 by Mildenhall et al. [43], NeRF is a form of neural rendering, a subset of generative AI, that generates novel views of a scene based on a partial set of 2D images. It achieves this by learning a mapping from 3D spatial coordinates and view directions $(x,y,z,\theta,\phi)$ to colors and density $(r,g,b,\sigma)$. This implicit representation allows NeRF to handle complex scenes with varying geometry and appearance, resulting in highly realistic renderings that include accurate lighting, shadows, and reflections. More detail can be found in Section 3.5.2.
@@
Creation Text [10, 44, 4, 45]
audio/music [46, 47] [48, 49, 50]
Image [46, 51] [24, 52, 53, 54, 55, 51, 56, 57, 58]
Animation/video [59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69] [70, 71, 72, 73, 63, 67, 69, 74, 75]
3D/AR/VR [76] [77, 78, 79, 80] [80, 81, 82]
Information Analysis Text categorization [83, 84, 85, 86]
Film analysis [87, 88, 89]
Content retrieval [90, 91, 92, 93, 94, 95] [96]
Intelligent assistants [97]
Content Enhancement [98, 99, 100, 101, 102, 103] [104, 105, 106, 107] [108]
Enhancement Style transfer [109, 110, 111] [112, 113] [110, 114]
and Post Super-resolution [115, 116, 117, 118, 119, 120, 103, 121, 122] [123, 124, 125, 126, 122] [127, 123, 128, 125, 129]
Production Restoration [130, 131, 119, 132, 103, 133, 115, 18, 134, 5, 135, 136, 137, 138, 139, 140, 141, 142, 141, 143] [106, 128, 144, 145, 146, 126, 147] [148]
Inpainting [149, 150, 151, 152, 153] [124, 128]
Fusion [154, 155, 156, 157] [158]
Editing/VFX [159] [159, 160]
Information Segmentation [161, 162, 163, 5, 164, 165, 166, 167, 168] [169, 170, 171] [172, 173]
Extraction Recognition [174, 11, 175, 16, 176, 17, 177, 166, 178, 179, 180] [181, 182, 183, 184]
and Tracking [185, 186, 187, 188, 189, 190, 191, 192, 193] [194, 195, 196] [197]
Understanding 3D Reconstruction [198, 199, 200, 201, 166, 202, 203] [204, 205, 206, 207] [43, 208, 209, 204, 210, 211, 212, 213, 214, 215, 216, 217, 203],[218, 219, 220, 221, 222, 223, 224, 225]
Compression Image [226, 227, 228] [229, 230, 231, 232] [233, 234, 235, 236]
Video [237, 238] [239] [240, 241, 40, 242, 243, 244, 245, 246, 247]
@@
Quality Image [248, 249, 250]
Assessment Video [251, 252, 253, 254, 255]
@@ 3 Advanced AI for the creative industries
Similarly to our previous (2021) review of AI for the creative industries [8], Table 1 categorizes applications and corresponding AI-based solutions. These areas are explored in more detail below.
@@ 3.1 Content creation
Content creation is a fundamental activity of artists and designers and the term ‘AI art’ refers to artforms created with the assistance of an AI algorithm or entirely by an AI system. This can refer to various digital forms including images, texts, audio, and videos. The roots of AI art can be traced back to the 20th century, exemplified by AARON, a computer program initiated in 1972 to autonomously produce paintings and drawings [256]. The practicality of AI art has been enhanced with advancements in deep learning, particularly GANs from 2014 and, more recently, transformers, DMs and INRs.
@@ 3.1.1 Text generation, script and journalism
In the era of LLMs, AI writing tools have been widely used to assist various writing tasks, including generation of written articles, blog posts, essays, and reports. These tools go beyond mere grammar and spelling checks; they boast advancements enabling them to analyze the style and tone of written material, adding images, videos and tables, offering suggestions to enhance clarity, coherence, and overall readability [257]. Moreover, AI tools extend their utility beyond content generation by automating tasks like keyword generation, meta tags, and descriptions, thereby increasing search rankings using search engine optimization (SEO). Additionally, they support the process of publishing across multiple online platforms. Transformers have been used to generate image captions by combining information from the images with a word prefix or questions [44].
@@
AI script generators serve as beneficial aids for writers, filmmakers, and game developers, offering inspiration, idea generation, and assistance in crafting entire scripts [2, 258]. Human-AI brainstorming is helpful and saves time [259]. Presently, there are numerous software and websites providing both free and paid script generation services. However, many of these tools are still constrained when it comes to longform creative writing. Dramatron, developed by Google [260], introduces hierarchical language generation, enabling the creation of cohesive scripts and screenplays spanning long ranges. This includes elements such as titles, characters, story beats, location descriptions, and dialogue.
@@
As discussed earlier, chatbots are now powered by LLMs, effectively simulating human conversation. These fundamental LLMs are specialized for specific tasks. For instance, journalist AI and blog AI writers [^12] For example, see https://tryjournalist.com/ generate content with layouts suitable for print or online publication. Additionally, AI tools exist that are designed to detect AI-generated content (e.g., for checking for copyright), AI-writing styles, content originality, and to ensure the naturalness and flow of articles. Undoubtedly, generative AI is reshaping the way artists and journalists operate. For an in-depth exploration of the impact and implications of these technological advancements on news organizations, refer to the survey conducted by Beckett et al. [261].
@@
Generating text and scripts automatically can also be done through image and video inputs without text prompts (e.g., image captioning [262]) and with text prompts. These approaches are referred to as Vision Language Models (VLMs): multimodal models that learn from images and text. The most common and prominent models often consist of an image encoder, an embedding projector to align image and text representations, often via a dense neural network, and a text decoder stacked in this order. The most well-known technique is Contrastive Language-Image Pre-training (CLIP) [263]. More recent work in [45] scales up the vision vocabulary by incorporating new image features into the existing CLIP model, resulting in improved content understanding. A comprehensive survey of VLMs for vision tasks can be found in [264].

### v5

@@ 2.3 Diffusion Models
A generative model, in the context of AI, exploits machine learning to learn a probability distribution of the training data to generate new data samples. The very first models were based on Autoencoders (AEs) that learn to encode input data into a lower-dimensional representation (latent space) and then decode it back to its original form. A specific type of AE, a variational autoencoders (VAE) [Kingma:auto:2014], learns the latent space as statistical parameters of probabilistic distributions, leading to significant improvement of the generated results. Concurrently, Goodfellow et al. [Goodfellow:GAN:2014] introduced an alternative architecture known as a Generative Adversarial Network (GAN). GANs comprise two competing AI modules: a generator, which creates a sample, and a discriminator, which determines whether the received sample is real or generated. When comparing VAEs to GANs, VAEs exhibit greater stability during training, whereas GANs excel at producing realistic images. More details about AEs and GANs for creative technologies can be found in our previous review [Anantrasirichai:AI:2022].
@@
An important factor in driving the rapid growth of generative AI has been the development of diffusion probabilistic models (referred to as diffusion models (DMs)). The first DM was introduced in 2015 by Sohl-Dickstein et al. [Dickstein:Deep:2015], using Nonequilibrium Thermodynamics. However, it took a further 5 years for DMs to generate desirable results: the era of DMs began with Denoising Diffusion Probabilistic Models (DDPMs) proposed by Ho et al. [Ho:DDPM:2020] in 2020 and Score-based diffusion models proposed by Song et al. [Song:Score:2021] in 2021. These involve a simplified process using a denoising autoencoder to approximate Bayesian inference. In brief, the models leverage a diffusion process to learn a probability distribution of the input data. As the name suggests, the data is diffused by gradually adding noise at each iteration step as shown in Fig. 1 (b). A deep neural network (DNN) is then trained to remove this noise, called the denoising process or reverse process. Consequently, the trained model uses random noise to generate data with characteristics similar to those of the training samples. Comparing to GANs, DMs provide higher diversity samples [Dhariwal:Diffusion:2021] and a training process that is much more stable and does not suffer from mode collapse. DMs are however computationally intensive and require longer training times compared to GANs. The complexity can be significantly reduced by training the DMs in latent space. Latent Diffusion Models (LDM) [Rombach:LDM:2022] use pretrained networks to convert images to feature maps, and perform training on a low-dimensional space. The diagram of LDM is shown in Fig. 1 (c).
@@
Generating a synthesized sample at random might not be particularly useful, especially for creative industry applications. Therefore, conditional diffusion models have been proposed, supporting a wide range of applications such as text-to-sound, text-to-images, and image-to-videos. For DMs, the conditional distributions are modeled using a conditional denoising autoencoder. Classifier guidance was introduced in [Dhariwal:Diffusion:2021] to improve the generation of images of a desired class. For example, when we provide the model with information, such as ‘a flower’, the DM will synthesize a variety of flower images, as the word ‘flower’ guides the model toward the latent distribution that is formed by various images of flowers. The work in [Choi:ILVR:2021] simply refines the latent space of well-trained unconditional DDPM so that the higher-level semantics of the synthetic samples are similar to the reference (conditioning). The LDM [Rombach:LDM:2022] offers more flexible conditional image generators by adding cross-attention layers (referred to Transformers in Section 2.1) to the denoising autoencoder. A survey on the methods and applications of DMs prior to 2024 can be found in [Cao:survey:2024].
@@ 2.4 Implicit Neural Representations
Implicit Neural Representations (INR), also called neural fields, neural implicits or coordinate-based neural networks, represent input content implicitly through learned functions $F$, as shown in Eq. 3. They can be considered as fields $x$ (represented by a scalar, vector, or a tensor with a value, such as a magnetic field in physics) that are fully or partially parameterized by a neural network $\Phi$, typically an MLP [xie2022neural].
@@
Although this concept appears complex, the process is actually very straightforward. For example, in the case of an image, the coordinates of each pixel $(x,y)$ contain color information $(r,g,b)$. The INR inputs $(x,y)$ to the MLP and learns to provide the output $(r,g,b)$. The weights and biases of the MLP now represent such an image. Usually, the number of parameters of the MLP is smaller than the total number of pixels multiplied by 3, accounting for the 3 color channels. Hence, one of its emerging applications is in data compression [kwan2024hinerv]. Moreover, the INR can handle complex and high-dimensional data efficiently, attracting attention for visual computing applications such as 3D scene reconstruction.
Traditional MLPs employ ReLU (rectified linear unit) for non-linear activation due to its simplicity. However, Sitzmann et al. [sitzmann:siren:2020] demonstrated that periodic functions, such as sinusoids, are more suitable for representing complex natural signals, offering a better fit to the first- and second-order derivatives of the signals. However, this activation can cause ringing artifacts. Saragadam et al. instead proposed using complex Gabor wavelets [Saragadam:wire:2023], which learn to represent high frequencies better and simultaneously are robust to noise.
One of the fastest-growing areas that exploits INRs is Neural Radiance Fields (NeRF), evidenced by 57 papers presented at CVPR, the largest annual conference in computer vision, in 2022 growing to 175 papers in 2023101010https://markboss.me/post/nerf_at_cvpr23/, before dropping to 71 in 2024, largely due to competition from 3D Gaussian Splatting [^11] https://github.com/Yubel426/NeRF-3DGS-at-CVPR-2024. First introduced in 2020 by Mildenhall et al. [Mildenhall:NeRF:2020], NeRF is a form of neural rendering, a subset of generative AI, that generates novel views of a scene based on a partial set of 2D images. It achieves this by learning a mapping from 3D spatial coordinates and view directions $(x,y,z,\theta,\phi)$ to colors and density $(r,g,b,\sigma)$. This implicit representation allows NeRF to handle complex scenes with varying geometry and appearance, resulting in highly realistic renderings that include accurate lighting, shadows, and reflections. More detail can be found in Section 3.5.2.
@@
Creation Text [Vaswani:attention:2017, Wang:SIMVLM:2022, openai:gpt4:2023, wei2024vary]
audio/music [Alayrac:Flamingo:2022, Huang:GenerSpeech:2022] [Li:Diffusion:2022, Yang:Diffsound:2023, Evans:stable:2025]
Image [Alayrac:Flamingo:2022, esser:scaling:2024] [Rombach:LDM:2022, Brooks:InstructPix2Pix:2023, gal:Image:2023, Gandikota:Unified:2024, Lian:LLMG:2024, esser:scaling:2024, ren:hypersd:2024, Feng_Ma_2025, Liu_Ma_2025]
Animation/video [hong:cogvideo:2023, villegas:phenaki:2023, Azadi:Make:2023, Yu:Bidirectionally:2023, Liu:FETV:2023, wang:disco:2024, xu:VASA-1:2024, corona:vlogger:2024, Gupta:Photorealistic:2024, Hu_2024_CVPR, Zhu:INFP:2024] [singer:Make:2023, molad:dreamix:2023, wang:modelscope:2023, wu:tune:2023, Liu:FETV:2023, Gupta:Photorealistic:2024, Zhu:INFP:2024, wang2025lavie, wu2025customcrafter]
3D/AR/VR [yang:Holodeck:2024] [Xu:NeuralLift:2023, Melas:RealFusion:2023, Qian:Magic123:2024, tang:dreamgaussian:2024] [tang:dreamgaussian:2024, ren:dreamgaussian4d:2023, zhao2024clear]
Information Analysis Text categorization [sun:text:2023, shi:chatgraph:2023, Hou:promptboosting:2023, AI2025125952]
Film analysis [Mao:Biases:2023, krugmann:sentiment:2024, Hartmann:More:2023]
Content retrieval [Metzler:Rethinking:2021, Yan:Universal:2023, Lu:content:2023, Rajput:recommender:2023, li:unigen:2024, li2024learning] [Jin:DiffusionRet:2023]
Intelligent assistants [King:Sasha:2024]
Content Enhancement [Xu:SNR:2022, liang:RVRT:2022, Wang:Ultra:2023, Lin:SPATIO:2024, Youk:FMA:2024, Liang:VRT:2024] [HOU:Global:2023, Yi:Diff:2023, Jiang:Low:2023, lin2024lowlight] [Yang:Implicit:2023]
Enhancement Style transfer [Deng:StyTr2:2022, Moon:generalizable:2023, Chung_2024_CVPR] [Zhang:Inversion:2023, Chai:StableVideo:2023] [Moon:generalizable:2023, Kim:Controllable:2024]
and Post Super-resolution [Liang:SwinIR:2021, Lu:Transformer:2022, Liu:Learning:2022, Chen:Activating:2023, li:GRL:2023, kang:gigagan:2023, Liang:VRT:2024, xu:videogigagan:2024, wang2025seedvr] [Saharia:image:2023, Moliner:solving:2023, Gao:Implicit:2023, cao2025zero, wang2025seedvr] [Chen:Learning:2021, Saharia:image:2023, Fei:Generative:2023, Gao:Implicit:2023, Yin:CLE:2023]
Production Restoration [Wang:Uformer:2022, Zamir:Restormer:2022, li:GRL:2023, yang:ldp:2023, Liang:VRT:2024, Morris:DaBiT:2024, Liang:SwinIR:2021, Fan:SUNet:2022, Yu:DBT:2022, Wang:Painter:2023, Song:vision:2023, Xu:Video:2023, mao:single:2022, Zhang:Image:2024, zou2024deturb, fang2025guided, Yue:RViDeformer:2025, Jin:Masked:2025, Yue:RViDeformer:2025, Shi:VmambaIR:2025] [Jiang:Low:2023, Fei:Generative:2023, yang:realworld:2023, Nair:AT-DDPM:2023, Jaiswal:Physics:2023, cao2025zero, feng2025residual] [Jiang:NeRT:2023]
Inpainting [Li:MAT:2022, Liu:Reduce:2022, Ren:DLFormer:2022, Zhou:ProPainter:2023, HUANG:Sparse:2024] [Moliner:solving:2023, Fei:Generative:2023]
Fusion [Ma:SwinFusion:2022, Rao:TGFuse:2023, Liu:Multi:2023, LI2024102147] [Zhao:DDFM:2023]
Editing/VFX [Shi:Motion-I2V:2024] [Shi:Motion-I2V:2024, guo2024liveportrait]
Information Segmentation [Bowen:Marked:2022, Kirillov:SAM:2023, Ke:SAM-HQ:2023, Wang:Painter:2023, Wang:SegGPT:2023, Zou:Segment:2023, Oquab:DINOv2:2024, ravi2024sam2, Zhang:DVISp:2025] [Wu:DiffuMask:2023, Xu:Open:2023, Gu:Diffusioninst:2024] [Gong:Continuous:2023, Cen:Segment:2023]
Extraction Recognition [Carion:DERT:2020, Dosovitskiy:image:2021, Zhu:Deformable:2021, Liu:Swin:2021, Neimark:video:2021, Liu:Swinv2:2022, Huang:MonoDTR:2022, Oquab:DINOv2:2024, zhao:videoprism:2024, im2025gate3d, tian2025yolov12] [Li:Your:2023, Chen:DiffusionDet:2023, Zhang:DiffAD:2025, WU2025102965]
and Tracking [Meinhardt:TrackFormer:2022, zeng:motr:2022, cui:mixformer:2022, Mayer:Transforming:2022, yang:track:2023, Chen:SeqTrack:2023, Zhang:MOTRv2:2023, Yi:Comprehensive:2024, kang2025exploring] [Luo:DiffusionTrack:2024, Xie:DiffusionTrack:2024, Zhang:DiffusionTracker:2024] [Jung:AnyFlow:2023]
Understanding 3D Reconstruction [Wang:multi:2021, Zhang:Lite:2023, Chen:Vision:2023, Yang:depthanything:2024, Oquab:DINOv2:2024, Yang:depthanythingv2:2024, LIU:DSEM:2025] [Barron:Mip-NeRF360:2022, Ji:DDP:2023, wynn:diffusionerf:2023, Ke:Repurposing:2024] [Mildenhall:NeRF:2020, pumarola:DNeRF:2020, mueller:instant:2022, Barron:Mip-NeRF360:2022, Mildenhall:NeRFDark:2022, Fang:Fast:2022, Guo:neural:2022, 9879447, Liu_2024_CVPR, azzarelli:waveplanes:2023, zhan2024kfd, Tang_2024_CVPR, LIU:DSEM:2025],[Fridovich:kplanes:2023, kerbl:3Dgaussians:2023, wu:4dgaussians:2024, Yu:CoGS:2024, Huang:SCGS:2024, Wang2025, junkawitsch2025eva, kong2025efficient]
Compression Image [zhu2022transformer, zou2022devil, liu2023learned] [careil2023towards, yang2024lossy, hoogeboom2023high, ghouse2023residual] [sitzmann2020implicit, dupont2021coin, dupontcoin++, strumpler2022implicit]
Video [xiang2022mimt, mentzer2022vct] [li2024extreme] [chen2021nerv, bai2023ps, kwan2024hinerv, kim2024c3, leguay2024cool, kwan2024nvrc, gao2024pnvc, ruan2024point, kwan2024immersive]
@@
Quality Image [cheon2021perceptual, golestaneh2022no, shi2024transformer]
Assessment Video [wu2022fast, feng2024rankdvqa, wu2023exploringvideo, he2024cover, peng2024rmt]
@@ 3 Advanced AI for the creative industries
Similarly to our previous (2021) review of AI for the creative industries [Anantrasirichai:AI:2022], Table 1 categorizes applications and corresponding AI-based solutions. These areas are explored in more detail below.
@@ 3.1 Content creation
Content creation is a fundamental activity of artists and designers and the term ‘AI art’ refers to artforms created with the assistance of an AI algorithm or entirely by an AI system. This can refer to various digital forms including images, texts, audio, and videos. The roots of AI art can be traced back to the 20th century, exemplified by AARON, a computer program initiated in 1972 to autonomously produce paintings and drawings [encyclopedia_ai_v1]. The practicality of AI art has been enhanced with advancements in deep learning, particularly GANs from 2014 and, more recently, transformers, DMs and INRs.
@@ 3.1.1 Text generation, script and journalism
In the era of LLMs, AI writing tools have been widely used to assist various writing tasks, including generation of written articles, blog posts, essays, and reports. These tools go beyond mere grammar and spelling checks; they boast advancements enabling them to analyze the style and tone of written material, adding images, videos and tables, offering suggestions to enhance clarity, coherence, and overall readability [ippolito:creative:2022]. Moreover, AI tools extend their utility beyond content generation by automating tasks like keyword generation, meta tags, and descriptions, thereby increasing search rankings using search engine optimization (SEO). Additionally, they support the process of publishing across multiple online platforms. Transformers have been used to generate image captions by combining information from the images with a word prefix or questions [Wang:SIMVLM:2022].
@@
AI script generators serve as beneficial aids for writers, filmmakers, and game developers, offering inspiration, idea generation, and assistance in crafting entire scripts [Jeary2024, Azzarelli:Reviewing:2024]. Human-AI brainstorming is helpful and saves time [guo:exploring:2024]. Presently, there are numerous software and websites providing both free and paid script generation services. However, many of these tools are still constrained when it comes to longform creative writing. Dramatron, developed by Google [Mirowski:cowriting:2023], introduces hierarchical language generation, enabling the creation of cohesive scripts and screenplays spanning long ranges. This includes elements such as titles, characters, story beats, location descriptions, and dialogue.
@@
As discussed earlier, chatbots are now powered by LLMs, effectively simulating human conversation. These fundamental LLMs are specialized for specific tasks. For instance, journalist AI and blog AI writers [^12] For example, see https://tryjournalist.com/ generate content with layouts suitable for print or online publication. Additionally, AI tools exist that are designed to detect AI-generated content (e.g., for checking for copyright), AI-writing styles, content originality, and to ensure the naturalness and flow of articles. Undoubtedly, generative AI is reshaping the way artists and journalists operate. For an in-depth exploration of the impact and implications of these technological advancements on news organizations, refer to the survey conducted by Beckett et al. [Beckett:Generating:2023].
@@
Generating text and scripts automatically can also be done through image and video inputs without text prompts (e.g., image captioning [Stefanini:From:2023]) and with text prompts. These approaches are referred to as Vision Language Models (VLMs): multimodal models that learn from images and text. The most common and prominent models often consist of an image encoder, an embedding projector to align image and text representations, often via a dense neural network, and a text decoder stacked in this order. The most well-known technique is Contrastive Language-Image Pre-training (CLIP) [radford2021learning]. More recent work in [wei2024vary] scales up the vision vocabulary by incorporating new image features into the existing CLIP model, resulting in improved content understanding. A comprehensive survey of VLMs for vision tasks can be found in [Zhang:vision:2024].